dielectric sphere of radius R is hollowed out in the region 0 ≤ r ≤ s and a

thin, grounded, conducting shell is inserted at r = s. The sphere is placed in a uniform,

external E⃗ -field E⃗ = E0ẑalong the z axis. The dielectric constant of the hollowed sphere is εr

. Calculate the potential in the region r ≥ R.


1
Expert's answer
2021-08-10T17:56:33-0400

Answer:-

We have given data :-

region as "0\le r\le S"

r=S

and "\overrightarrow{\in}=\in_o \overrightarrow{z}"

"\in_r" = Dielectric Constant



Assume , before hollowed out a Sphere (R1) the charge on whole sphere is 'Q'

And after hollowed out , and charged to q [ Dielectric filled in the cavity ]

"\bull" Relation between q and Q

"\boxed{q=Q(1-\frac{1}{\in_r})}" "\rightarrow" this is charge left on hollow Spherical Shell.

"\bull" Electrical Field due to'Q' at a distance 'r'

"|\in (r)|=k.\frac{Q}{r^2}" "[\because k=\frac{1}{4\pi\in_o}]"

"\in_o =k\frac{Q}{r^2}\\
\boxed{Q=\frac{1}{k}\in_o.r^2}"


therefore , "q=\frac{\in_o}{k}r(1-\frac{1}{\in_r})"

Electric potential (V) due to charge (q) at a distance 'r' given by

"V=k.\frac{q}{r}\\
\ \ \ \ = k.\frac{\in_or}{k}(1-\frac{1}{e_r})\\"

"\therefore \boxed{V=\in_or(1-\frac{1}{e_r})}"